OpenStudy (anonymous):

CALCULUS 3: What are the possible angles between 2 unit vectors if e x f is 1/2?

OpenStudy (anonymous):

pi/6 and 5pi/6?

OpenStudy (anonymous):

\[e\times f=||e||~||f||\sin\theta\] Since \(e\) and \(f\) are unit vectors, \[\frac{1}{2}=\sin\theta\] So yes, \(\dfrac{\pi}{6}\) and \(\dfrac{5\pi}{6}\) work.